OpenStudy (anonymous):

trigonometry: Two legs of a triangle have a length 5 and 7, and the angle they determine has measure 2pi/3 radians. Determine the area of the triangle.

OpenStudy (anonymous):

area = (1/2) *5*7*sin(2pi/3 )=(35*sqrt(3))/4

OpenStudy (anonymous):

appx 15.156

OpenStudy (anonymous):

thank you @matricked and @sai1234 !

OpenStudy (anonymous):

area is 1/2 *a *bsintheta =0.5 *5 *7 *sin2pi/8 =0.5 *5 *7 *1/root(2) =12.3

OpenStudy (anonymous):

@jessicajuanng mention not